# CoherenDream: Boosting Holistic Text Coherence in 3D Generation via Multimodal Large Language Models Feedback

### v1

@@ Abstract
Score Distillation Sampling (SDS) has achieved remarkable success in text-to-3D content generation. However, SDS-based methods struggle to maintain semantic fidelity for user prompts, particularly when involving multiple objects with intricate interactions. While existing approaches often address 3D consistency through multiview diffusion model fine-tuning on 3D datasets, this strategy inadvertently exacerbates text-3D alignment degradation. The limitation stems from SDS’s inherent accumulation of view-independent biases during optimization, which progressively diverges from the ideal text alignment direction. To alleviate this limitation, we propose a novel SDS objective, dubbed as Textual Coherent Score Distillation (TCSD), which integrates alignment feedback from multimodal large language models (MLLMs). Our TCSD leverages cross-modal understanding capabilities of MLLMs to assess and guide the text-3D correspondence during the optimization. We further develop 3DLLaVA-CRITIC - a fine-tuned MLLM specialized for evaluating multiview text alignment in 3D generations. Additionally, we introduce an LLM-layout initialization that significantly accelerates optimization convergence through semantic-aware spatial configuration. Comprehensive evaluations demonstrate that our framework, CoherenDream, establishes state-of-the-art performance in text-aligned 3D generation across multiple benchmarks, including T3Bench and TIFA subset. Qualitative results showcase the superior performance of CoherenDream in preserving textual consistency and semantic interactions. As the first study to incorporate MLLMs into SDS optimization, we also conduct extensive ablation studies to explore optimal MLLM adaptations for 3D generation tasks.
Figure 1: While original Score Distillation suffer from per-view bias accumulation resulting in objects omission, Textual Coherent Score Distillation first incorporates feedback of MLLM into SDS optimization, leading to textual consistent 3D results.
1 Introduction
@@
3D content generation is essential for diverse applications, including gaming, virtual reality, and robotics simulation. Recently, significant progress has been made in text-to-3D generation through Score Distillation Sampling (SDS) [30, 41, 21, 4, 34, 26]. SDS-based methods enable high-quality and diverse 3D generation based on user-provided text inputs, effectively distilling image distributions from 2D diffusion models [32] into parameterized 3D representations.
@@
Despite these advances, significant challenges persist in generating 3D content that faithfully adheres to user prompts, especially those involving multiple objects and intricate interactions. The primary limitation stems from SDS’s per-view distillation mechanism, which lacks explicit constraints for maintaining global textual consistency across multiple updates. Per-view independent optimization accumulates biases during the optimization process, causing a progressive deviation from the intended text-to-3D alignment. Furthermore, while recent variants of SDS [34, 24, 18] adopt fine-tuned diffusion models on specific 3D datasets to enhance 3D consistency, they exacerbate the problem of textual inconsistency in 3D generations [15, 18], leading to object omissions and physically implausible spatial relationships, as illustrated in Fig. 1.
@@
How to develop coherent text-to-3D generation remains relatively unexplored. Recent research efforts, such as JointDreamer [15], emphasize the importance of maintaining textual consistency within the original diffusion model compared to its variants [34]. DreamView [44] focuses on training a view-specific text-to-diffusion model to enhance alignment with specific viewpoints. However, both approaches still face the challenge of bias accumulation during SDS optimization. Another intuitive way to alleviate textual inconsistency in 3D generation is through compositional strategies, such as GALA3D [46] and GraphDreamer [9]. Nevertheless, GALA3D [46] suffers from non-overlap decomposition, which leads to unnatural spatial relationships. Although GraphDreamer [9] advances the modeling of object relationships through scene graphs, it remains limited by compositional optimization, resulting in implausible object combination. In contrast, our work aims to generate coherent results within a holistic 3D representation where object relationships and spatial arrangements maintain textual consistency across all viewpoints.
@@
Recent advances in multimodal large language models (MLLMs) demonstrate powerful cross-modal understanding capabilities, leading to successful integrations in text-to-image generation  [37, 12, 8, 19, 5]. While 2D counterparts are promising, directly adapting MLLMs to 3D generation is not feasible, owing to their limited grasp of 3D representation and spatial relationships. We address these limitations by reformulating text-3D alignment as across-view question-answering tasks. Our key insight is that MLLMs’ semantic reasoning capabilities can complement gradient updates from image distribution when properly contextualized within SDS optimization. To this end, we present CoherenDream, which firstly regards MLLMs as dynamic semantic assessor to ensure textual consistency in SDS optimization. We encode the input text and the across-view image distribution from the diffusion model as text-format ground truth, measuring the loss against the text sequences predicted by the MLLM to serve as feedback. This introduces Textual Coherent Score Distillation (TCSD), which adopts MLLM feedback in the SDS optimization to steer the optimization toward a textual-consistent distribution.
@@
Current MLLMs are primarily designed as language assistants, they lack the proficiency to critique 3D generation effectively. To bridge this gap, we further develop a 3DLLaVA-CRITIC based on LLaVA-OneVision 0.5b [17] to enhance the quality of the feedback. Specifically, we design a view-aware data collection pipeline to simulate the gradient updates during SDS optimization, ensuring the delivery of accurate feedback. Furthermore, we introduce LLM-layout initialization, which is integrated with TCSD to warm up the 3D representation, thereby enhancing the textual consistency for image distribution from diffusion models. Extensive experimental results demonstrate that CoherenDream not only generates text-coherent 3D content but also outperforms other text-to-3D generation methods in terms of quality and quantitative metrics.
@@ 2 Related Works
The Score Distillation Sampling (SDS) algorithm has achieved surprising results in text-to-3D generation. It pioneers by [30], utilizing 2D diffusion model priors [32] to optimize 3D representations. Recent advancements have further refined this technique by enhancing 3D representations [21, 4, 45, 39], improving generation quality [14, 41, 47, 20], and ensuring 3D consistency [15, 34, 18, 33, 2]. Despite these impressive results, these methods still struggle with multi-object prompts [10] and semantic interaction. These challenges often stem from the lack of global textual consistency and accumulation of view-independent biases. To address it, we introduce MLLM feedback into SDS, which dynamically revises the update direction of 3D representations, guiding the score distillation process toward text-3D alignment.
Figure 2: The overview of CoherenDream framework. CoherenDream involves LLM-Layout initialization, semantic coherent score distillation and 3DLLaVA-CRITIC with three kinds of guidance tasks, producing text-3D aligned results from MLLM feedback.
@@
One intuitive approach for text-aligned 3D generation is decomposing holistic representations into individual components for separate optimization before recombination. However, existing methods [22, 3] suffer from quality degradation due to the challenges in managing layout constraints during NeRF optimization and their dependence on potentially inaccurate predefined layouts. Recently, GALA3D [46] attempts to use 3D Gaussians representation while dynamically refine layout during generation. However, its non-overlapping constraint leads to unnatural spatial relationships and scale inconsistencies (e.g., excessive spacing between the girl and dinosaur in Fig. 4). GraphDreamer [9] advances the field by incorporating automatically parsed scene graphs to guide object interactions. Nevertheless, it inherits the fundamental limitations of compositional approaches, struggling to maintain holistic coherence, particularly in scenes requiring intimate object interactions (e.g., faceless cat in Fig. 4). In this work, we initialize holistic 3D representation through LLM-layout, which serves as a better warm-up for SDS optimization. The proposed TCSD further enhance optimization towards text-3D alignment. Our approach can generate holistic 3D content while facilitating more realistic interactions among various objects.
@@
Recent advancements in large language models (LLMs)[40, 29, 1] have led to increased interest in Multimodal Large Language Models (MLLMs), which combine vision understanding capabilities with traditional LLMs[29, 17]. Given their promising abilities in visual reasoning and understanding, MLLMs are being explored to enhance 3D generation, particularly in areas such as automatic evaluation [42, 10] and data preparation [38, 6]. However, existing approaches have not yet integrated MLLMs directly into the 3D generation process. In contrast, our work fine-tunes the LLaVA model [17] to provide semantic coherence criticism and view-checking based on multi-view images. We are the first to leverage MLLM guidance to assist directly in SDS optimization.
3 Preliminaries
@@
Score Distillation Sampling (SDS) employs priors from pre-trained 2D diffusion models to facilitate the generation of 3D content, which is widely adopted in text-to-3D methods [30, 21, 4, 41, 25]. Given a user input prompt $y$, parameterized 3D representation $\theta$, and pre-trained 2D diffusion model $\Phi(\textbf{x}_{t}|y)$ along with noise prediction network $\epsilon_{\Phi}(\textbf{x}_{t},t,y)$, the objective of SDS is to optimize $\theta$ by minimizing the KL-divergence between the rendered image distribution $q_{t}^{\theta}(\textbf{x}_{t}|c)$ and sampled image distribution $p_{t}(\textbf{x}_{t}|y))$ from diffusion model:
@@
Ignoring the UNet Jacobian [30], the gradient computation of SDS loss is as follows:
$\small\begin{split}\nabla_{\theta}\mathcal{L}_{SDS}(\theta)&\triangleq\mathrm{E}_{t,\textbf{x}}[w(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}D_{KL}(q_{t}^{\theta}(\textbf{x}_{t}|c,y)||p_{t}(\textbf{x}_{t}|y))]\\ &\triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\epsilon_{\Phi}(\textbf{x}_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\end{split}$ (2)
@@
4 Method
In this section, we introduce CoherenDream, a novel text-aligned 3D generation framework as depicted in Fig. 2. We first demonstrate the derivation of Textual Coherent Score Distillation (TCSD) in Sec. 4.1, which introduce MLLM feedback for across-view image distribution. Then we introduce a fine-tuned MLLM, 3DLLaVA-CRITIC to better evaluate across-view alignment in 3D generation in Sec. 4.2. Finally, we elaborate on the overall framework CoherenDream in Sec. 4.3, where we integrate three guidance tasks with TCSD and novel LLM-layout initialization technique to further enhance textual consistency.
4.1 Textual Coherent Score Distillation
In accordance with Eq. 1, we observe significant limitations in textual understanding and reasoning within the original SDS framework: (1) $p_{t}(\textbf{x}_{t}|y)$ is constrained by diffusion models, which may diverge considerably from the actual user-prompt distributions [16, 37]; (2) $p_{t}(\textbf{x}_{t}|y)$ heavily relies on the rendered image x of last updated 3D representation $\theta$, leading to bias accumulation. To address these limitations, we introduce informative feedback from MLLMs to $p_{t}(\textbf{x}_{t}|y)$ towards textual consistent distribution.
We first formulate the ideal textual consistent distribution $\hat{p}_{t}(\textbf{x}_{t}|y)$ as a joint distribution $\hat{p}_{t}(\textbf{x}_{t}|y)=p_{t}(\textbf{x}_{t}|y)p(y|\textbf{x}_{t})$. Based on the connection between diffusion models and score matching [36], the score function of $p_{t}(\textbf{x}_{t}|y)$ can be derived as:
@@ 4.1 Textual Coherent Score Distillation
$\small\nabla_{\textbf{x}_{t}}\text{log}(p_{t}(\textbf{x}_{t}|y)p(y|\textbf{x}_{t}))=-\frac{1}{\sqrt{1-\alpha_{t}}}\epsilon_{\Phi}(\textbf{x},t,y)+\nabla_{\textbf{x}_{t}}\text{log}\,p(y|\textbf{x}_{t})$ (4)
@@
$\begin{split}\small\hat{\epsilon}_{\Phi}(\textbf{x}_{t},t,y)=\epsilon_{\Phi}(\textbf{x}_{t},t,y)-\sqrt{1-\alpha_{t}}\nabla_{\textbf{x}_{t}}\text{log}\,p(y|\textbf{x}_{t})\\ =\epsilon_{\Phi}(\textbf{x}_{t},t,y)+\lambda\cdot\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\textbf{x}_{t}))\end{split}$ (5)
@@
$\begin{split}\small\delta(\textbf{T},\hat{\textbf{x}}_{0})&=\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\hat{\textbf{x}}_{0}))\\ &=\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\frac{\textbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\Phi}(\textbf{x}_{t},t,y)}{\sqrt{\alpha}}))\end{split}\vspace{-3mm}$ (6)
@@
$\begin{split}\small\nabla_{\theta}\mathcal{L}_{TCSD}(\theta)\triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\hat{\epsilon}_{\Phi}(\textbf{x}_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\\ \triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\epsilon_{\Phi}(\textbf{x}_{t},t,y)+\lambda\cdot\delta(\textbf{T},\hat{\textbf{x}}_{0})-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\end{split}$ (7)
4.2 3DLLaVA-CRITIC
Original MLLMs are primarily designed as language assistants and cannot assess 3D generation adequately. To bridge this gap, we decompose the evaluation of textual consistency in 3D generation into three question-answering tasks: scene graph generation, multi-label image classification, and view classification. Building on this foundation, we introduce 3DLLaVA-CRITIC, following the architecture of LLaVA-OV [17] and fine-tuned using instruct-tuning pairs generated by advanced LLMs such as GPT-4o [29].
4.2.1 Task definition.
@@
Figure 3: View-aware data collection pipeline for 3DLLaVA-CRITIC that consist of (1) using LLM to generate diverse coarse text prompt and corresponding layout and rendering from random viewpoints in Blender; (2) random sampling images from T2I diffusion model conditioned on layout image and randomly assembling into a 2$\times$2 grid image; (3) employing GPT-4o to extract semantic annotation, including scene graph and view direction.
4.2.2 View-aware data collection pipeline.
@@
Unlike natural images, the sampled image distribution in SDS is conditioned on noisy rendered images from randomly sampled camera poses. To achieve better assessment during SDS optimization, we introduce a view-aware data collection pipeline, as illustrated in Fig. 3. Detailed prompts for each step in the pipeline are available in the Appendix. Layout condition generation. Inspired by LayoutGPT [7], we utilize GPT-4 [29] to generate 200 diverse text prompts based on examples in the DreamFusion [30] library. Layout is defined as 3D bounding boxes with box center coordinates and dimension: (x, y, z, h, w, l). These layouts are rendered in  [^1] https://www.blender.org/Blender, and we randomly sample camera poses to produce 32 rendering images per layout.
@@
View-aware image generation. Based on the previously created prompts and layout condition images, we enhance the original prompts by integrating view prompts (e.g., side, back, front, overhead). Utilizing DeepFloyd-IF [35] and MVDream [34], we generate view-aware images by varying random seeds, noise strengths, and denoising steps, in line with the optimization process of CoherenDream. Under each layout condition, we randomly select 4 camera views to create a 2$\times$2 grid image, resulting in a total of 36 grid images.
@@
GPT-4o Annotation. After the last step, we curate 307,409 grid images and then prompt GPT-4o [29] to generate scene graphs, view prompts, and perform object extraction. These high-quality grid image-text pairs serve as instruction tuning data for proposed 3DLLaVA-CRITIC.
4.3 Framework of CoherenDream
@@
Building upon the proposed TCSD optimization, we present the CoherenDream framework. This framework is based on NeRF [27] as 3D representation, utilizing Instant-NGP [28] with a volume renderer. To ensure coherence with user input, we incorporate three guidance tasks (as outlined in Sec. 4.2.1) with Eq. 7, which direct sampled image distribution toward the textual consistent distribution. During the optimization process, we employ well-established techniques such as time-annealing [14] and resolution scaling-up [41]. In addition, we introduce a novel technique called LLM-layout initialization, which significantly enhances the quality of the generated 3D content.
@@
Guidance Tasks. We equipe TCSD optimization with three feedbacks from 3DLLaVA-CRITIC: in scene graph generation, multi-label classification and view classification. Detailed task descriptions can be found in Sec. 4.2.1. For the scene graph generation task, $\textbf{T}_{sg}$ can be extracted by humans or GPT according to user’s input. And answers $\textbf{T}_{ob}$ of multi-label classification can be directly obtained from $\textbf{T}_{sg}$. $\textbf{T}_{view}$ can be decided given the sampled camera pose. Therefore, the overall feedback of CoherenDream is demonstrated as: $\delta_{\text{CoherenDream}}=\lambda_{sg}\delta(\textbf{T}_{sg},\hat{\textbf{x}}_{0})+\lambda_{ob}\delta(\textbf{T}_{ob},\hat{\textbf{x}}_{0})+\lambda_{view}\delta(\textbf{T}_{view},\hat{\textbf{x}}_{0})$.

### v3

@@ Abstract
Score Distillation Sampling (SDS) has achieved remarkable success in text-to-3D content generation. However, SDS-based methods struggle to maintain semantic fidelity for user prompts, particularly when involving multiple objects with intricate interactions. While existing approaches often address 3D consistency through multiview diffusion model fine-tuning on 3D datasets, this strategy inadvertently exacerbates text-3D alignment degradation. The limitation stems from SDS’s inherent accumulation of view-independent biases during optimization, which progressively diverges from the ideal text alignment direction. To alleviate this limitation, we propose a novel SDS objective, dubbed as Textual Coherent Score Distillation (TCSD), which integrates alignment feedback from multimodal large language models (MLLMs). Our TCSD leverages cross-modal understanding capabilities of MLLMs to assess and guide the text-3D correspondence during the optimization. We further develop 3DLLaVA-CRITIC - a fine-tuned MLLM specialized for evaluating multiview text alignment in 3D generations. Additionally, we introduce an LLM-layout initialization that significantly accelerates optimization convergence through semantic-aware spatial configuration. Our framework, CoherenDream, achieves consistent improvement across multiple metrics on TIFA subset.As the first study to incorporate MLLMs into SDS optimization, we also conduct extensive ablation studies to explore optimal MLLM adaptations for 3D generation tasks.
Figure 1: Textual Coherent 3D Generation with CoherenDream. By integrating multimodal LLM feedback into SDS optimization, our Textual Coherent Score Distillation corrects view-bias drift and yields faithful, text-aligned 3D content.
Introduction
@@
3D content generation is essential for diverse applications, including gaming, virtual reality, and robotics simulation. Recently, significant progress has been made in text-to-3D generation through Score Distillation Sampling (SDS) (Poole et al. 2023; Wang et al. 2024; Lin et al. 2023a; Chen et al. 2023; Shi et al. 2024; Metzer et al. 2023). SDS-based methods enable high-quality and diverse 3D generation based on user-provided text inputs, effectively distilling image distributions from 2D diffusion models (Rombach et al. 2022) into parameterized 3D representations.
@@
Despite these advances, significant challenges persist in generating 3D content that adheres to user prompts, especially those involving multiple objects and intricate interactions. The limitation stems from SDS’s per-view distillation, which lacks global consistency constraints. Independent per-view optimization accumulates biases and progressively deviates from the intended text-to-3D alignment. Furthermore, while recent variants of SDS (Shi et al. 2024; Liu et al. 2024; Li et al. 2024b) adopt fine-tuned diffusion models on specific 3D datasets to enhance 3D consistency, they exacerbate the problem of textual inconsistency in 3D generations (Jiang et al. 2025; Li et al. 2024b), leading to object omissions and physically implausible spatial relationships.
@@
How to develop coherent text-to-3D generation remains relatively unexplored. Recent research efforts, such as JointDreamer (Jiang et al. 2025), emphasize the importance of maintaining textual consistency within the original diffusion model compared to its variants (Shi et al. 2024). DreamView (Yan et al. 2025) trains a view-specific text-to-diffusion model to enhance alignment with specific viewpoints. However, both approaches still face the challenge of bias accumulation during SDS optimization. Another intuitive way to alleviate textual inconsistency in 3D generation is through compositional strategies, such as GALA3D (Zhou et al. 2024) and GraphDreamer (Gao et al. 2024). Nevertheless, GALA3D (Zhou et al. 2024) suffers from non-overlap decomposition, which leads to unnatural spatial relationships. Although GraphDreamer (Gao et al. 2024) advances the modeling of object relationships through scene graphs, it remains limited by compositional optimization, resulting in implausible object combination. In contrast, our work aims to generate coherent results within a holistic 3D representation where object relationships and spatial arrangements maintain textual consistency across all viewpoints.
@@
Recent advances in multimodal large language models (MLLMs) demonstrate powerful cross-modal understanding capabilities, leading to successful integrations in text-to-image generation  (Sun et al. 2023; Hu et al. 2024; Feng et al. 2024b; Lian et al. 2023; Cho, Zala, and Bansal 2023). While 2D counterparts are promising, directly adapting MLLMs to 3D generation is not feasible, owing to their limited grasp of 3D representation and spatial relationships. We address these limitations by reformulating text-3D alignment as across-view question-answering tasks. Our key insight is that MLLMs’ semantic reasoning capabilities can complement gradient updates from image distribution when properly contextualized within SDS optimization, as shown in Fig 1. To this end, we present CoherenDream, which firstly regards MLLMs as dynamic semantic assessor to ensure textual consistency in SDS optimization. We encode the input text and the across-view image distribution from the diffusion model as text-format ground truth, measuring the loss against the text sequences predicted by the MLLM to serve as feedback. This introduces Textual Coherent Score Distillation (TCSD), which adopts MLLM feedback in the SDS optimization to steer the optimization toward a textual-consistent distribution.
@@
Current MLLMs are primarily designed as language assistants, they lack the proficiency to critique 3D generation effectively. To bridge this gap, we further develop a 3DLLaVA-CRITIC based on LLaVA-OneVision 0.5b (Li et al. 2024a) to enhance the quality of the feedback. Specifically, we design a view-aware data collection pipeline to simulate the gradient updates during SDS optimization, ensuring the delivery of accurate feedback. Furthermore, we introduce LLM-layout initialization, which is integrated with TCSD to warm up the 3D representation, thereby enhancing the textual consistency for image distribution from diffusion models. Extensive experimental results demonstrate that CoherenDream not only generates text-coherent 3D content but also outperforms other text-to-3D generation methods in terms of quality and quantitative metrics.
@@
The Score Distillation Sampling (SDS) algorithm has achieved surprising results in text-to-3D generation. It pioneers by (Poole et al. 2023), utilizing 2D diffusion model priors (Rombach et al. 2022) to optimize 3D representations. Recent advancements have further refined this technique by enhancing 3D representations (Lin et al. 2023a; Chen et al. 2023; Yi et al. 2024; Tang et al. 2024), improving generation quality (Huang et al. 2023; Wang et al. 2024; Zhu, Zhuang, and Koyejo 2024; Liang et al. 2024), and ensuring 3D consistency (Jiang et al. 2025; Shi et al. 2024; Li et al. 2024b; Seo et al. 2024; Armandpour et al. 2024). Despite these impressive results, these methods still struggle with multi-object prompts (He et al. 2023) and semantic interaction. These challenges often stem from the lack of global textual consistency and accumulation of view-independent biases. To address it, we introduce MLLM feedback into SDS, which dynamically revises the update direction of 3D representations, guiding the score distillation process toward text-3D alignment.
@@
An intuitive approach for text-aligned 3D generation is to decompose holistic representations into individual components for separate optimization before recombination. However, existing methods (Lin et al. 2023b; Bai et al. 2023) suffer from quality degradation due to the challenges in managing layout constraints during NeRF optimization and their dependence on potentially inaccurate predefined layouts. Recently, GALA3D (Zhou et al. 2024) attempts to use 3D Gaussians representation while dynamically refine layout during generation. However, its non-overlapping constraint leads to unnatural spatial relationships and scale inconsistencies. GraphDreamer (Gao et al. 2024) advances the field by incorporating automatically parsed scene graphs to guide object interactions. Nevertheless, it inherits the fundamental limitations of compositional approaches, struggling to maintain holistic coherence, particularly in scenes requiring intimate object interactions. In this work, we warm-up a holistic 3D representation through LLM-layout for SDS optimization. The proposed TCSD further enhance optimization towards text-3D alignment. Our approach can generate holistic 3D content while facilitating more realistic interactions among various objects.
@@
Recent advancements in large language models (LLMs)(Touvron et al. 2023; OpenAI 2023; Anil et al. 2023) have led to increased interest in Multimodal Large Language Models (MLLMs), which combine vision understanding capabilities with LLMs(OpenAI 2023; Li et al. 2024a). Given their promising abilities in visual reasoning and understanding, MLLMs are being explored to enhance 3D generation, particularly in areas such as automatic evaluation (Wu et al. 2024b; He et al. 2023) and data preparation (Sun et al. 2024; Fang et al. 2024). However, existing approaches have not yet integrated MLLMs directly into the 3D generation process. In contrast, our work fine-tunes the LLaVA model (Li et al. 2024a) to provide semantic coherence criticism and view-checking based on multi-view images. We are the first to use MLLM guidance to assist directly in SDS optimization.
Preliminaries
@@
Score Distillation Sampling (SDS) employs priors from pre-trained 2D diffusion models to facilitate the generation of 3D content, which is widely adopted in text-to-3D methods (Poole et al. 2023; Lin et al. 2023a; Chen et al. 2023; Wang et al. 2024; Luo et al. 2024). Given a user input prompt $y$, parameterized 3D representation $\theta$, and pre-trained 2D diffusion model $\Phi(\textbf{x}_{t}|y)$ along with noise prediction network $\epsilon_{\Phi}(\textbf{x}_{t},t,y)$, the objective of SDS is to optimize $\theta$ by minimizing the KL-divergence between the rendered image distribution $q_{t}^{\theta}(\textbf{x}_{t}|c)$ and sampled image distribution $p_{t}(\textbf{x}_{t}|y))$ from diffusion model:
@@
Ignoring the UNet Jacobian (Poole et al. 2023), the gradient computation of SDS loss is as follows:
$\begin{split}\nabla_{\theta}\mathcal{L}_{SDS}(\theta)&\triangleq\mathrm{E}_{t,\textbf{x}}[w(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}D_{KL}(q_{t}^{\theta}(\textbf{x}_{t}|c,y)||p_{t}(\textbf{x}_{t}|y))]\\ &\triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\epsilon_{\Phi}(\textbf{x}_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\end{split}$ (2)
@@
Figure 2: Overview of CoherenDream framework. CoherenDream involves LLM-Layout initialization, textual coherent score distillation and 3DLLaVA-CRITIC with three kinds of guidance tasks, producing text-3D aligned results from MLLM feedback.
Method
In this section, we introduce CoherenDream, a novel text-aligned 3D generation framework as depicted in Fig. 2. We first demonstrate the derivation of Textual Coherent Score Distillation (TCSD), which introduce MLLM feedback for across-view image distribution. Then we introduce a fine-tuned MLLM, 3DLLaVA-CRITIC to better evaluate across-view alignment in 3D generation. Finally, we elaborate on the overall framework CoherenDream, where we integrate three guidance tasks with TCSD and novel LLM-layout initialization technique to further enhance textual consistency.
Textual Coherent Score Distillation (TCSD)
In accordance with Eq. 1, we observe significant limitations in textual understanding and reasoning within the original SDS framework: (1) $p_{t}(\textbf{x}_{t}|y)$ is constrained by diffusion models, which may diverge considerably from the actual user-prompt distributions (Kirstain et al. 2023; Sun et al. 2023); (2) $p_{t}(\textbf{x}_{t}|y)$ heavily relies on the rendered image x of last updated 3D representation $\theta$, leading to bias accumulation. To address these limitations, we introduce informative feedback from MLLMs to $p_{t}(\textbf{x}_{t}|y)$ towards textual consistent distribution.
We first formulate the ideal textual consistent distribution $\hat{p}_{t}(\textbf{x}_{t}|y)$ as a joint distribution $\hat{p}_{t}(\textbf{x}_{t}|y)=p_{t}(\textbf{x}_{t}|y)p(y|\textbf{x}_{t})$. Based on the connection between diffusion models and score matching (Song and Ermon 2019), the score function of $p_{t}(\textbf{x}_{t}|y)$ can be derived as:
@@
$\nabla_{\textbf{x}_{t}}\text{log}(p_{t}(\textbf{x}_{t}|y)p(y|\textbf{x}_{t}))=-\frac{1}{\sqrt{1-\alpha_{t}}}\epsilon_{\Phi}(\textbf{x},t,y)+\nabla_{\textbf{x}_{t}}\text{log}\,p(y|\textbf{x}_{t})$ (4)
@@
$\begin{split}\hat{\epsilon}_{\Phi}(\textbf{x}_{t},t,y)=\epsilon_{\Phi}(\textbf{x}_{t},t,y)-\sqrt{1-\alpha_{t}}\nabla_{\textbf{x}_{t}}\text{log}\,p(y|\textbf{x}_{t})\\ =\epsilon_{\Phi}(\textbf{x}_{t},t,y)+\lambda\cdot\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\textbf{x}_{t}))\end{split}$ (5)
@@
$\begin{split}\delta(\textbf{T},\hat{\textbf{x}}_{0})&=\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\hat{\textbf{x}}_{0}))\\ &=\nabla_{\textbf{x}_{t}}\mathcal{L}_{ce}(\textbf{T},f_{cr}(\frac{\textbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\Phi}(\textbf{x}_{t},t,y)}{\sqrt{\alpha}}))\end{split}$ (6)
@@
$\begin{split}\nabla_{\theta}\mathcal{L}_{TCSD}(\theta)\triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\hat{\epsilon}_{\Phi}(\textbf{x}_{t},t,y)-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\\ \triangleq\mathrm{E}_{t,\epsilon_{\Phi}}[w(t)(\epsilon_{\Phi}(\textbf{x}_{t},t,y)+\lambda\cdot\delta(\textbf{T},\hat{\textbf{x}}_{0})-\epsilon)\frac{\partial g(\theta,c)}{\partial\theta}],\end{split}$ (7)
3DLLaVA-CRITIC
Original MLLMs are primarily designed as language assistants and cannot assess 3D generation adequately. To bridge this gap, we decompose the evaluation of textual consistency in 3D generation into three question-answering tasks: scene graph generation, multi-label image classification, and view classification. Building on this foundation, we introduce 3DLLaVA-CRITIC, following the architecture of LLaVA-OV (Li et al. 2024a) and fine-tuned using instruct-tuning pairs generated by GPT-4o (OpenAI 2023). Detailed prompts can be found in the the Appendix.
Task Definition
@@
View-aware Data Collection Pipeline
@@
Unlike natural images, the sampled image distribution in SDS is conditioned on noisy rendered images from randomly sampled camera poses. To achieve better assessment during SDS optimization, we introduce a view-aware data collection pipeline, as illustrated in Fig. 3. Detailed prompts for each step in the pipeline are available in the Appendix.
Layout condition generation. Inspired by LayoutGPT (Feng et al. 2024a), we utilize GPT-4 (OpenAI 2023) to generate 200 diverse text prompts based on examples in the DreamFusion (Poole et al. 2023) library. Layout is defined as 3D bounding boxes with box center coordinates and dimension: (x, y, z, h, w, l). These layouts are rendered in Blender, and we randomly sample camera poses to produce 32 rendering images per layout.
@@
View-aware image generation. Based on the previously created prompts and layout condition images, we enhance the original prompts by integrating view prompts (e.g., side, back, front, overhead). Utilizing DeepFloyd-IF (Shonenkov et al. 2023) and MVDream (Shi et al. 2024), we generate view-aware images by varying random seeds, noise strengths, and denoising steps, in line with the optimization process of CoherenDream. Under each layout condition, we randomly select 4 camera views to create a $\times2$ grid image, resulting in a total of 36 grid images.
@@
GPT-4o Annotation. After the last step, we curate 307,409 grid images and then prompt GPT-4o (OpenAI 2023) to generate scene graphs, view prompts, and perform object extraction. These high-quality grid image-text pairs serve as instruction tuning data for proposed 3DLLaVA-CRITIC.
Figure 3: View-aware data collection pipeline for 3DLLaVA-CRITIC that consist of (1) using LLM to generate diverse coarse text prompt and corresponding layout and rendering from random viewpoints in Blender; (2) random sampling images from T2I diffusion model conditioned on layout image and randomly assembling into a $\times2$ grid image; (3) employing GPT-4o to extract semantic annotation, including scene graph and view direction.
Framework of CoherenDream
@@
Building upon TCSD optimization, we present the CoherenDream framework that is based on NeRF (Mildenhall et al. 2021) as 3D representation, utilizing Instant-NGP (Müller et al. 2022) with a volume renderer. To ensure coherence with user input, we incorporate three guidance tasks with Eq. 7, which direct sampled image distribution toward the textual consistent distribution. During the optimization process, we employ well-established techniques such as time-annealing (Huang et al. 2023) and resolution scaling-up (Wang et al. 2024). In addition, we introduce a novel technique called LLM-layout initialization, which significantly enhances the quality of the generated 3D content.
Guidance Tasks.
@@
We equipe TCSD optimization with three feedbacks from 3DLLaVA-CRITIC: in scene graph generation, multi-label classification and view classification. Detailed task descriptions can be found before. For the scene graph generation task, $\textbf{T}_{sg}$ can be extracted by humans or GPT according to user’s input. And answers $\textbf{T}_{ob}$ of multi-label classification can be directly obtained from $\textbf{T}_{sg}$. $\textbf{T}_{view}$ can be decided given the sampled camera pose. Therefore, the overall feedback of CoherenDream is demonstrated as: $\delta_{\text{CoherenDream}}=\lambda_{sg}\delta(\textbf{T}_{sg},\hat{\textbf{x}}_{0})+\lambda_{ob}\delta(\textbf{T}_{ob},\hat{\textbf{x}}_{0})+\lambda_{view}\delta(\textbf{T}_{view},\hat{\textbf{x}}_{0})$.